# Resolving the Exploitation-Exploration Dilemma in Evolutionary Algorithms: A Novel Human-Centered Framework

### v1

@@ Abstract
Evolutionary Algorithms (EAs) are powerful tools for tackling complex computational problems, yet effectively managing the exploitation and exploration dynamics—crucial for robust search navigation—remains a persistent challenge for EA designers, and leads to the so-called exploration-exploitation dilemma. In this paper, a new human-centered framework is proposed to resolve this dilemma. Unlike the traditional approach, the search process will not be compromised of a single-phase nor the decision-maker tuning efforts will be distributed among the algorithm’s traditional parameters such as defining new evolutionary operators internal to the algorithm to influence its search navigation. Instead, a human-centered two-phase search process, compromised of a global search phase followed by a local phase will be utilized. In this framework, the designer plays the central role in directing the algorithm’s search navigation through the focused tuning efforts of a new Search Space Size Control parameter external to the algorithm which proves itself to be the dominant parameter in-effect to the algorithm’s effective search navigation.
The framework is applicable to any search algorithm. We demonstrate its effectiveness on 14 well-known benchmark problems in unconstrained optimization.
Keywords Exploration-Exploitation Dilemma  $\cdot$ Evolutionary Algorithms  $\cdot$ Search Space Size Control Parameter  $\cdot$ Human-Centered Two-Phase Search Framework  $\cdot$ Global Optimization
@@ 1 Introduction
EAs are a family of population-based, stochastic adaptation models that simulate mechanisms and patterns observed in evolutionary biology and the natural ecosystem. The central idea in these algorithms is to allow a population of artificial creatures—e.g., strings that represent candidate solutions to a problem or programs that perform a given task—to build themselves up through an artificial evolutionary process that aims to approximate what mother-nature does. Over the years, various types of EAs have been proposed, including Evolution Strategies (ES)[1], Evolutionary Programming (EP), Genetic Algorithms (GAs) [2], and Genetic Programming (GP) [3]. The key differences between these algorithms lie in the type of representation used and how self-adaptation is approximated through search operators [4]
By design, EAs are particularly well-suited to solving optimization problems, especially those involving complex, nonlinear, or dynamic environments. This realization can be traced back to the pioneering work of computer scientists in the 1950s and 1960s, who first explored the potential of evolutionary processes as optimization tools for engineering problems [5], and since then, EAs have evolved into a powerful class of computational techniques, capable of addressing diverse problems across science, engineering, and industry.
A system undergoing a natural evolutionary process must continuously compute decisions to carry itself from its current state to a more fit state. The computation of these decisions relies on two complementary sub-processes: exploration and exploitation. Exploration involves the search for resources that may enhance the system chances of survival, and improvement while, exploitation focuses on optimizing the use of known resources to maximize immediate benefits/state-transition. For example, in biological systems, exploration can manifest as genetic mutations, behavioral changes, or the colonization of new habitats, all of which contribute to the genetic diversity and resilience of a population ([6], [7]), while exploitation is essential for refining existing traits or behaviors of a population that have proven successful, thereby enhancing the overall system’s fitness and performance in its current environment.
The interplay between exploration and exploitation is crucial for the long-term success of any evolutionary system. A system that leans too heavily on exploration may become inefficient, wasting resources on unproductive paths, while one that overly focuses on exploitation may become stagnant, failing to adapt to new challenges ([8], [9]). Thus, an effective evolutionary strategy requires a dynamic balance between these two processes to allow the system to adaptively respond to both immediate needs and future uncertainties ([2], [10]).
The same requirement holds for a robust EA. However, unlike nature, which intelligently manages this dilemma, machines do not achieve it inherently. Therefore, human algorithm designers must carefully craft mechanisms that enable EAs to explore the solution space effectively while exploiting promising solutions to ensure convergence, efficiency and optimization.
In the context of algorithm design, the exploitation-exploration dilemma comes down to the following question: How could the algorithm’s capacity for exploration be maximized without compromising its ability to exploit promising solutions? On one hand, there is a desire for the algorithm to thoroughly explore the feasible solution space to ensure comprehensive coverage and the potential discovery of optimal solutions. On the other hand, it is equally crucial to prevent an overamplification of randomness within the algorithm, as this can impede convergence and degrade overall performance.
@@
A review of the literature indicates that the current approach practiced in the literature for addressing this dilemma in EAs primarily centers around fine-tuning and adjusting traditional parameters within EAs. These traditional parameters include: the selection of evolutionary operators, operator probabilities, population size, and others. In this context, researchers often engage in the development of new operators internal to the algorithm and experiment with other various parameter configurations in pursuit of optimizing algorithm performance. And indeed, while specific tuning configurations can lead to improvements, like the selection of particular EA operators outperforming others in specific contexts, the progress achieved through this approach typically exhibits an incremental nature. Furthermore, we argue that this approach inherently faces limitations in delivering effective, intelligible, and noticeable advancements in addressing the dilemma in question.
This paper proposes a novel human-centered two-phase search framework (HCTPS) to resolve this dilemma, and in which, exploration is maximized, but not at the expense of the exploitation potential of the algorithm. The framework consists of two distinct phases: the global search phase, in which the algorithm distributes its internal exploitation-exploration potential across the entire search cube $I$, followed by a local search phase, in which a sequence of selected sub-cubes $\{I_{i}\}$ from the original search cube $I$ (where $i=1,2,\dots,n$) undergoes further intensive sequential exploration. The rigorous implementation of the HCTPS is supported by the introduction of the human-centered search space control parameter (HSSCP), which allows the decision-maker to dynamically specify and adjust the sequence of sub-cubes to explore in the local search phase.
To demonstrate the effectiveness of HCTPS, we integrate the it with the canonical GA and evaluate its performance on 14 well-known benchmark unconstrained optimization problems. The results demonstrate the framework’s ability to enhance exploration while preserving the GA’s inherent robust exploitation, offering a scalable and generalizable solution for various search algorithms.
The remainder of this paper is designed to be as self-contained as possible, with an emphasize on clarity and the use of analogies for intuitive and comprehensive explanation: the next section presents a comprehensive overview of the canonical GA, including its key operators, terminology, and parameters. Section (3) illustrates the inherent limitations of traditional approach in managing the exploration-exploitation trade-off. Section (4) introduces the proposed HCTPS framework, detailing its conceptual design, the two-phase search process, and the role of the human-centered search space control parameter (HSSCP) in enabling dynamic and adaptive exploration, supported by a relatable analogy for intuitive accessibility. Section (5) evaluates the effectiveness of the HCTPS framework through experimental results obtained from benchmark unconstrained optimization problems, comparing its performance with the traditional approach, and ends with a reflective discussion, while section (6) concludes the paper and outlines potential directions for future research.
@@ 2 The Canonical Genetic Algorithm (GA)
GAs are among the most important and widely applied EAs for solving complex computational tasks ([11], [12]). The first variant of GAs was introduced by Holland in the 1960s and later refined in the 1970s through collaborations at the University of Michigan. As the name suggests, GAs draw inspiration from Darwinian principles of evolution, particularly natural selection, genetic variation, and the survival-of-the-fittest paradigm ([13]).
In this work, we adopt the function-optimizer interpretation of the GA and specifically implement Holland’s original formulation, commonly referred to in the literature as the simple or canonical GA. The progression of the GA is organized into six key steps, implemented through specific GA operators that collectively shape its evolutionary cycle. These steps are detailed below.
@@ 2.1 Terminology, Key Steps, and Operators
Alternatively, one may begin with a randomly generated subset of strings of length $L$ from a chosen representation system (e.g., binary, ternary, etc.) and then apply a DECODE function to map these strings back into the original feasible set $S$.
@@
GA designers acknowledge this trade-off by employing selection operators that balance the need for promoting fitness with maintaining population diversity. These operators are typically classified based on how they assign probabilities of selection—commonly referred to as selection pressure—to chromosomes within the population. Selection methods can be broadly divided into two categories: roulette-wheel selection and ordinal methods.
@@
The specific definition and implementation of the crossover operator can vary, and its effectiveness greatly depends on designing a suitable recombination mechanism. Various approaches exist, such as single-point crossover, multi-point crossover, uniform crossover, and other specialized methods tailored to the problem domain.
@@
There are many variations of the mutation operators. The common aspect shared by all the variations is the involvement of one or more changes that are made to the genes of the chromosomes. In other words, mutation performs a random walk in the vicinity of a candidate solution, adding a level of stochasticity and diversity to the GA.
@@
Finally comes the last step, after the recombination and mutation step which results in forming a new population, we need to introduce the new generated population in replacement with the old one through the replacement operator. There are many replacement operators to help achieve this task such as elitist replacement, generation-wise replacement, and steady-state replacement. These operators determine how the chromosomes from the new population are integrated into the next generation, ensuring the continuous evolution of the population over successive generations.
@@ 2.2 Parameters
A parameter in the context of an optimization algorithm is a configuration variable that influences the algorithm’s search navigation, behaviour, performance, and convergence properties. Formally, for an algorithm $\mathcal{A}$, operating on a solution space $\mathcal{X}$, a parameter $\theta\in\Theta$, where $\Theta$ is the parameter space, serves as an input to configure the algorithm but is not subject to optimization during its execution. This can be expressed as:
$\mathcal{A}:(\mathcal{X},\theta)\rightarrow\mathcal{Y},$
where $\theta$ adjusts the mapping between inputs $\mathcal{X}$ and outputs $\mathcal{Y}$.
Parameters can be classified based on their impact, update strategies, and temporal behaviour.
  • Temporal behaviour: Some parameters act as internal configuration variables and evolve during the optimization process, while others act as external configuration variables set before the optimization process, and remain static throughout the algorithm’s execution. Examples:
    • Internal configuration variable: The mutation probability $p_{m}$ in a GA can be set internally based on the algorithm’s feedback. If the algorithm is not converging well, the mutation rate can be increased to introduce more exploration in the search space.
    • External configuration variable: The population size is set externally by the practitioner before running the GA and remains constant throughout the optimization process.
  • Updating strategy: Parameters in an optimization process may either self-update based on feedback from the algorithm or be fixed by the practitioner. Examples:
    • Self-updated parameter: In a GA, $p_{m}$ can be dynamically adjusted based on the fitness improvement of the best solutions found. If the fitness improves rapidly, $p_{m}$ can be reduced to focus more on exploiting the promising solutions, or in a self-adaptive penalty-based GA where the penalty factor gets self-updated.
    • Practitioner-updated parameter: The maximum number of iterations, and the selection of the GA operators are typically specified by the practitioner and remains fixed throughout the optimization process.
  • Impact: Some parameters have more impact than others on the algorithm’s output/behavior in the sense that a small change in their value affects the algorithm’s output, even with slight adjustments, while others may require larger changes to produce noticeable effects.
Below, we analyze the traditional GA parameters in regards to their impact on the exploration-exploitation of the search navigation.
@@
    Population size Population size is a critical parameter for balancing the exploration and exploitation dynamics of a GA. A small population size limits exploration by reducing diversity, which increases the risk of premature convergence to suboptimal solutions ([14]). Conversely, a larger population size promotes exploration by sampling a broader range of the solution space, reducing the likelihood of stagnation in local optima. However, this comes at the cost of slower convergence, which may hinder the algorithm’s exploitation of promising solutions.
@@
    Crossover probability $p_{c}$. The crossover operator is the primary search mechanism GAs and one of their defining features compared to other EAs ([15]). The crossover probability $p_{c}$ controls the likelihood that two parent solutions will undergo crossover to produce offspring. A higher $p_{c}$ accelerates convergence by intensifying the recombination of genetic material, potentially leading to faster exploitation of promising solutions. However, excessive crossover may reduce population diversity, hindering the algorithm’s ability to explore the solution space effectively.
@@
    Selection of genetic operators. The choice of GA operatore—including selection, crossover, mutation, and replacement—determines the descriptive content (or the logic) of the mechanism by which individuals are chosen for reproduction, how genetic material is recombined, and how it is altered to produce new solutions.
@@
    Number of fitness function evaluations. Increasing the number of evaluations allows the algorithm to gather more information about the problem space, enhancing its ability to explore diverse regions and identify potential solutions in a meaningful way. At the same time, this provides opportunities to refine high-performing solutions, improving accuracy and solution quality through exploitation. However, excessive evaluations can significantly increase computational costs, slowing down the algorithm, while too few evaluations may result in premature convergence and suboptimal solutions.

### v2

@@ Abstract
Evolutionary Algorithms (EAs) are widely employed tools for complex search and optimization tasks; however, the absence of an overarching operational framework that permits a systematic regulation of the exploration-exploitation tradeoff—critical for efficient convergence—restricts the full actualization of their potential, leading to the so-called exploration-exploitation dilemma in algorithm design. A systematic resolution to this dilemma requires: (1) independent yet coordinated control over exploration and exploitation, and (2) an explicit, computationally feasible, adaptive regulation mechanism. The current, almost decentralized, traditional parameter tuning-centeric approach—lacks the foundation to satisfy these requirements under encoding-imposed structural constraints.
We propose a Human-Centered Two-Phase Search (HCTPS) framework, in which the actualization of (1) and (2) is enabled through an external configuration variable—the Search Space Control Parameter (SSCP). As the sole control knob of HCTPS, the SSCP centralizes exploration adjustments, sparing users from micromanaging traditional parameters with unintelligible interdependencies. To this construct, the human user serves as a meta-parameter, adaptively steering the regulatory process via SSCP adjustments.
We prove that the HCTPS strictly surpasses the current approach in terms of search space coverage without disrupting the EAs’ inherent convergence mechanisms, demonstrate a concrete instantiation of it—using the Genetic Algorithm as the underlying heuristic on a suite of global benchmark unconstrained optimization problems, provide a through assessment of the proposed framework, and envision future research directions. Any search algorithm prone to this dilemma can be applied in light of the proposed framework, being algorithm-agnostic by design.
Keywords Exploration-Exploitation Dilemma  $\cdot$ Evolutionary Algorithms  $\cdot$ Human-centered Computation  $\cdot$ Search Space Size Control Parameter  $\cdot$ Human-Centered Search Framework  $\cdot$ Nature-Inspired Algorithms
@@ 1 Introduction
Evolutionary Algorithms (EAs) are a family of stochastic computational models that approximate the processes of genetic variation (variation), selection pressure (natural selection), adaptation to selection pressure (adaptation), and reproductive heritability observed in biological evolution.
Variation refers to the differences in characteristics among organisms within a population. These differences naturally arise due to small changes in inherited traits, which can result from genetic mutations, recombination, or other natural processes. Natural selection refers to the survival challenges imposed by the environment, while adaptation is the process by which genetic variations that enhance an organism’s fitness in its respective environment become more prevalent in a population over generations through natural selection, and reproductive heritability ensures that beneficial adaptations are passed down to subsequent generations. Together, these processes contribute to steering biological systems toward progressively more robust and well-adapted states.
EAs enable systems of artificial entities—e.g., strings representing candidate solutions to a problem or programs executing specific tasks—to build themselves up over many iterations in response to a selection pressure exerted by a given fitness criterion. First, a random population of artificial entities is generated, and their performance is evaluated according to this fitness criterion which then rewards the fitter entities with a higher survival advantage, allowing them to reproduce through mutation and recombination. Over successive iterations, this evolution ideally yields entities with progressively higher fitness—e.g., solutions that more effectively solve a given problem or programs capable of executing the intended tasks with near-optimal precision—thus filling the available niche in the artificial ecosystem.
Over the years, numerous variants of EAs have been proposed. The primary distinctions between them lie in their representation schemes—how evolving entities are encoded—and the manner in which evolutionary processes are discretized and implemented. Notable examples include Evolution Strategies (ES)(Rechenberg, 1984), Evolutionary Programming (EP), Genetic Algorithms (GAs) (Holland, 1975), and Genetic Programming (GP) (Koza, 1992).
By design, EAs are particularly well-suited for optimization problems, particularly those characterized by complex, highly nonlinear, and multimodal function landscapes, and where traditional optimization algorithms often struggle. This realization can be traced back to the pioneering work of computer scientists in the 1950s and 1960s, who first explored the potential of evolutionary processes as optimization tools for engineering problems (Mitchell, 1998). While EAs are mostly applied to optimization problems, their applicability extends to a wide range of problem domains across science, engineering, artificial intelligence, arts, music, knowledge discovery, and more. Beasley (2000) overviews some of the application areas of EAs.
The course of adaptation in natural evolutionary systems is a result of maintaining the proper tradeoff between two implicit forces: exploration and exploitation. Exploration is induced by the diversification mechanisms that give rise to genetic variability within a population, while exploitation is induced by natural selection, which systematically capitalizes on genetic variations that enhance organisms’ fitness. An improper tradeoff between these two forces can destabilize systems, and if persistently maintained, may ultimately lead to a total collapse.
Unlike machines, nature is remarkably effective at self-regulating this exploration-exploitation tradeoff, evidenced by the continuation of life on Earth. Therefore, it is often very difficult to engineer artificial evolutionary systems (e.g. EAs) capable of regulating this tradeoff autonomously. This difficulty leads to the so-called exploration-exploitation dilemma. This dilemma, first introduced in the EAs literature by authors of (Eiben & Schippers, 1998), refers to the challenge of optimally allocating computational resources, in the proper ratio, to the exploration and exploitation mechanisms within the algorithm to ensure efficient evolution with respect to the problem landscape under consideration.
1.1 Requirements for a Systematic Resolution
A systematic resolution to this dilemma necessitates a principled framework in which the following are met:
  1. 1.
    Decoupled and independent control: Exploration and exploitation must be regulated through distinct and independently adjustable mechanisms while maintaining controlled coordination to prevent search instability. This decoupling ensures that adjustments to exploration do not implicitly degrade exploitation or vice versa, thereby preventing entangled tradeoffs that hinder search efficiency.
  2. 2.
    Explicit and adaptive regulation: The framework must provide an explicitly defined, dynamically adjustable exploration control mechanism that can be adapted in response to measurable search progress indicators (e.g., search space coverage, fitness diversity,final algorithm output).
  3. 3.
    Theoretical and computational feasibility: The framework must be algorithm-agnostic, and its internal mechanisms must remain computationally scalable without requiring an exponential increase in representation set size, prohibitively large population sizes, or exhaustive parameter tuning.
@@ 1.1 Requirements for a Systematic Resolution
To the best of our knowledge, the current approach practiced in the literature for addressing this dilemma in EAs primarily centers around distributing tuning efforts among traditional parameters, such as the selection of evolutionary operators, operator probabilities, population size, and others. In this context, one often engages in the development of new operators internal to the algorithm and experiment with other various parameter configurations in pursuit of optimizing algorithm search performance. And indeed, while specific tuning configurations can lead to improvements—such as certain EA operators outperforming others in specific contexts— this approach does not amount to a systematic resolution to the dilemma.
The unintelligible interdependencies among these parameters obscure their individual contributions to exploration and exploitation. This prevents independent, controlled modulation of the exploration-exploitation tradeoff, and does not provide a principled basis for adaptive control, thus prohibiting the actualization of requirements (1), and (2).
Moreover, this approach under encoding-induced structural constraints further restrict the algorithm’s ability to enhance exploration without excessive computational overhead, thus prohibiting the actualization of requirement (3). Hence, the need for a new framework that provides the foundations for systematic resolution of the dilemma in question.
1.2 Contribution
The primary contributions of this work are as follows:
  1. 1.
    Rigorous background for context: We provide background sections on the canonical GA and the canonical EA framework to offer a fresh, rigorous, and stand-alone overview of standard EA mechanics and parameters, which culminate in key insights that motivate and inform the proposed HCTPS framework.
  2. 2.
    Technical analysis of current approach limitations: We analyze how the unintelligible interdependencies among traditional EA parameters, coupled with encoding-based structural constraints, prevent principled, adaptive regulation of exploration and exploitation in details.
  3. 3.
    Novel Human-Centered Two-Phase Search (HCTPS) framework: We introduce a new framework that externalizes exploration from the EA’s internal mechanics, placing it under a single, user-driven configuration variable called the Search Space Control Parameter (SSCP). We show how this decoupling provides a systematic basis for resolving the exploration–exploitation dilemma.
  4. 4.
    Proof of strictly enhanced exploration: Through a rigorous argument, we prove that HCTPS strictly increases the search space coverage compared to the current approach for an arbitrary EA, ensuring that the exploration-exploitation tradeoff can be modulated without degrading convergence mechanisms.
  5. 5.
    Empirical validation: We demonstrate a concrete instantiation of the HCTPS framework—using canonical GA as the underlying heuristic into a two-phase, human-centered search process on a set of global unconstrained blackbox optimization problems and compare results against the current approach, and we comment on the output.
  6. 6.
    Assessment of the proposed framework: We evaluate HCTPS in terms of its ability to meet the earlier requirements, conduct a cost-benefit analysis, and discuss wider applicability.
1.3 Structure of the Paper
The remainder of this paper is structured to be as self-contained as possible, with an emphasis on clarity, and the use of analogies to facilitate intuitive and comprehensive explanations and is organized around the contributions identified above as follows:
Sections 23 deliver contribution 1, while section 4 contribution 2, and section 5 contributions 3 and 4. Section 6 delivers contribution 5, and section 7 contribution 6. Finally section 8 summarizes the paper, offers key insights, concludes the work, and presents future research directions.
@@ 2 The Canonical Genetic Algorithm (GA)
The first variant of GAs was introduced by Holland in the 1960s and later refined in the 1970s through collaborations at the University of Michigan. It is commonly referred to in the literature as the simple or canonical GA. Like all EAs, GAs are inspired by the Darwinian theory of evolution (Darwin, 1859) but their terminology is derived from natural genetics.
The progression of the GA is organized into six key steps, each implemented through specific stochastic operators that collectively shape its evolutionary cycle. These steps are detailed below.
@@ 2.1 Terminology, Key Steps, and Operators
Alternatively, one may begin with a randomly generated subset of strings of length $L$ from a chosen representation system (e.g., binary, ternary, etc.) and then apply a DECODE function, which should also be bijective, to map these strings back into the original feasible set $S$.
@@
GA designers acknowledge this trade-off by employing selection operators that balance the need for promoting fitness with maintaining population diversity. These operators are typically classified based on how they assign probabilities of selection—commonly referred to as selection pressure—to chromosomes within the population. Common selection operators include:
  1. 1.
    Roulette wheel selection (fitness-proportional selection): Assigns selection probabilities proportional to fitness values (Goldenberg, n.d.). While favoring fitter chromosomes, it can lead to premature convergence if fitness differences are extreme.
  2. 2.
    Ordinal selection (rank-based selection): Assigns selection probabilities based on rank rather than raw fitness values pressure (Baker, 1985). Instead of using fitness directly, individuals are ranked, and probabilities are distributed according to their position. This method ensures a more controlled selection pressure, preventing any single individual from overwhelming the population, thus maintaining genetic diversity.
@@
The specific definition and implementation of the crossover operator can vary, and its effectiveness greatly depends on designing a suitable recombination mechanism. Many variants of the crossover operator exist, such as single-point, multi-point, uniform crossover (detailed description of these can be found in (De Jong, 2006); geometric crossover (Michalewicz, 2013), heuristic crossover (Wright, 1991), and partially mapped crossover (PMX) (Goldberg & Lingle, 1985).
@@
There are many variations of the mutation operators such as bit-flipping mutation (for binary representation), uniform mutation, non-uniform mutation, Gaussian mutation and Cauchy mutation; a comprehensive overview of these operators can be found in (Liu & Han, 2003). The common aspect shared by all the variations is the involvement of one or more changes that are made to the genes of the chromosomes. In other words, mutation performs a random walk in the vicinity of a candidate solution, adding a level of stochasticity and diversity to the GA.
@@
The final step in the GA is replacement, where the algorithm determines how to incorporate newly generated chromosomes into the next generation. This process is carried out using a replacement operator, which plays a crucial role in shaping the diversity and convergence rate of the algorithm. There are several variants of the replacement operator, each influencing the evolutionary process differently. Among the most common are: elitist replacement, generation-wise replacement, and steady-state replacement; a comprehensive overview can be found in (Eiben & Smith, 2015).
@@ 2.2 Parameters
A parameter in the context of an algorithm is a configuration variable that influences the algorithm’s behaviour, performance, and convergence properties, whose value remains fixed during the algorithm’s execution.
Formal definition: Let $\mathcal{A}$ be an algorithm operating on a input space $\mathcal{X}$ and producing outputs in the space $\mathcal{Y}$, represented as:
$\mathcal{A}:\mathcal{X}\times\Theta\rightarrow\mathcal{Y}.$
Where $\theta$ is a parameter to $\mathcal{A}$ and $\Theta$ denotes the space of all possible parameter values, then $\theta$’s role is to define the conditions under which the mapping applies.
Parameters can be classified based on their temporal behavior, updating strategy, and impact on algorithm performance.
  • Temporal behaviour: Parameters’ values in an algorithm can either remain static throughout the execution of all the iterations or change dynamically based on the algorithm’s progress according to some rule.
    • Dynamic parameter: Some parameters are adjusted between iterates in response to feedback from the algorithm. For example, the mutation probability $p_{m}$ in a GA may be modified if the algorithm is not converging effectively—increasing to enhance exploration or decreasing to focus on exploitation.
    • Static parameter: Other parameters are defined before execution and remain unchanged throughout all the iterates. For instance, the population size in a GA is typically set before the algorithm starts running and stays constant throughout the whole run.
  • Updating strategy: Parameters during the algorithm’s run may either self-update based on feedback from the algorithm or be set and updated by the human user (human-centered). In the first, we say such parameter is mechanically-centered, and in the latter human-centered. Examples:
    • Mechanically-centered parameter: The mutation probability $p_{m}$ in a GA can be dynamically adjusted based on the fitness improvement of the best solution found. If the fitness improves rapidly, $p_{m}$ may be decreased to focus efforts on exploitation; otherwise, it may be increased to encourage exploration.
    • Human-centered parameter: These parameters remain fixed during execution. Examples include the maximum number of iterations and the selection of genetic operators (e.g., crossover and mutation type).
    • Impact on algorithm behavior: Parameters vary in their impact on the algorithm’s performance. Some parameters have a stronger influence, meaning even small changes can significantly alter the search behavior, while others require larger adjustments to produce noticeable effects.
    Remark 2.
    The parameter definition presented here applies not only to computational algorithms, which are a specific class of mappings or operators, but more generally to any operator or a system of operators. In the latter, parameters define the structural and operational constraints under which the system functions.
    Below, we analyze the traditional GA parameters in regards to their impact on the exploration-exploitation of the search space.
@@
      Population size Population size is a critical parameter for balancing the exploration and exploitation dynamics of a GA. A small population size limits exploration by reducing diversity, which increases the risk of premature convergence to sub-optimal solutions (Roeva et al., 2013). Conversely, a larger population size promotes exploration by sampling a broader range of the solution space, reducing the likelihood of stagnation in local optima. However, this comes at the cost of slower convergence, which may hinder the algorithm’s exploitation of promising solutions.
@@
      Crossover probability $p_{c}$. The crossover operator is the primary search mechanism GAs and one of their defining features compared to other EAs (Tian, 2001). The crossover probability $p_{c}$ controls the likelihood that two parent solutions will undergo crossover to produce offspring. A higher $p_{c}$ accelerates convergence by intensifying the recombination of genetic material, potentially leading to faster exploitation of promising solutions. However, excessive crossover may reduce population diversity, hindering the algorithm’s ability to explore the solution space effectively.
@@
      Selection of genetic operators. The choice of GA operator—including selection, crossover, mutation, and replacement—determines the descriptive content (or the logic) of the mechanism by which individuals are chosen for reproduction, how genetic material is recombined, and how it is altered to produce new solutions.
@@
      Number of fitness function evaluations. Increasing the number of evaluations allows the algorithm to gather more information about the problem space, enhancing its ability to explore diverse regions and identify potential solutions in a meaningful way. At the same time, this provides opportunities to refine high-performing solutions, improving accuracy and solution quality through exploitation. However, excessive evaluations can significantly increase computational costs, slowing down the algorithm, while too few evaluations may result in premature convergence and sub-optimal solutions.
@@
3 The Canonical Evolutionary Algorithm (EA)
So far we discussed in detail the conceptual framework and key steps of the canonical GA. However, this framework is not exclusive to GAs; it applies to all EAs, meaning that the same six-step structure in (2) applies to all EAs. In the introduction of this paper, we briefly mentioned several notable examples of EAs, and outlined that they differ in two key areas: Representation schemes and the approximation of evolutionary processes employed.
In the next subsection, we give a summary to how these differences manifest among various EAs, followed by key insights into EA design and broad implications.
3.1 Key Differences among EAs
EAs can be distinguished based on two fundamental design aspects:
  1. 1.
    Representation schemes:
    • Genetic algorithms (GA): Use bit-string encoding or real-valued representations.
    • Evolution strategies (ES): Employ real-valued vectors, making them more suited for continuous optimization.
    • Genetic programming (GP): Uses tree-based structures, enabling the evolution of symbolic expressions or programs.
  2. 2.
    Approximation of evolutionary processes:
    • Selection mechanisms: GAs often rely on fitness-proportionate or tournament selection, whereas ES employs deterministic selection.
    • Reproduction mechanisms: GAs primarily use crossover and bitwise mutations, ES applies Gaussian perturbations, and GP modifies tree structures via subtree recombination.
    • Replacement mechanisms: GAs typically replace the entire population per generation, while ES and GP often incorporate elitist or steady-state replacement strategies to maintain diversity.